OpenStudy (anonymous):

You watched 280 people play a game of coin toss at a fair. Thirty-five people won. What is the experimental probability of winning the game?

OpenStudy (ranga):

35 people out of 280 people won the game. What is the probability of winning the game?

OpenStudy (ranga):

Probability of winning = Number of people who won / Number of people who tried = 35 / 280 = (5 * 7) / (5 * 7 * 8) = 1 / 8.